OpenStudy (anonymous):

FIND A^ IF VECTOR A =I^-J^+2K^

OpenStudy (unklerhaukus):

A = (x, -y, 2z)

OpenStudy (anonymous):

what^ , thats wrong!

OpenStudy (anonymous):

A^ I assume means A with a hat over the top. That means they want a vector of unit length The length of the vector = sqrt{1^2+1^2+2^2}= sqrt(6)

OpenStudy (anonymous):

\[( \frac{1}{\sqrt{6}}, -\frac{1}{\sqrt{6}}, \frac{2}{\sqrt{6}})\]

OpenStudy (unklerhaukus):

your quite right elecengineer, the hats (carets) confused me

OpenStudy (unklerhaukus):

A^ \[= A / |A| = 1/\sqrt{6} \times (x,-y,2z)\]